A is the angle of the incline when a block of mass m just starts slipping down. The distance covered by the block if thrown up the incline with an initial speed g is ?Ans: v^2/4gsinA

Grade:11

1 Answers

Arun
6 years ago
 

 

 

block moves down the plane with constant velocity it means net force is zero...

 

force due to friction (f) = mgsin A                              (A is the angle of plane)

 

net retardation force when body is projected upward is ( f+ mgsin A)

 

 f = mgsin A so

 

net retardation force = 2mgsin A

 

     ma = mgsin A

 

     a(retardation) = 2gsin A        .............1

 

 V2 = U2 - 2aS

 

finally velocity becomes 0 so

 

  V2 = 2aS

 

 S = V2/2a

 

    =V2/4gsin A